OpenStudy (anonymous):

Find the derivative of each of the the following functions. Show your work. f(x)=√((7+(√x3)) g(x)=7x/(√(5−2x)) Please help. I can't find the derivative of these 2 questions.

OpenStudy (anonymous):

i can't read them

OpenStudy (anonymous):

is it \[\large f(x)=\sqrt{7+\sqrt{x^3}}\]?

OpenStudy (anonymous):

yes and the second one is \[7x \div \sqrt{5-2x}\]

OpenStudy (anonymous):

\[\frac{d}{dx}[\sqrt{f(x)}]=\frac{f'(x)}{2\sqrt{f(x)}}\] by the chain rule

OpenStudy (anonymous):

if you put \(f(x)=7+\sqrt{x^3}=7+x^{\frac{3}{2}}\) then \(f'(x)=\frac{3}{2}\sqrt{x}\)

OpenStudy (anonymous):

shouldn't it be f(x)=\[1/(2\sqrt{7+x^3/2})\] ?